Butch will miss an important TV program while taking his statistics exam, so he sets both his VCRs to record it. The first one records 70% of the time, and the second one records 60% of the time. What is the probability that he gets home after the exam?


1
Expert's answer
2021-05-31T13:13:58-0400

Total problem:

Butch will miss an important TV program while taking his statistics

exam, so he sets both his VCRs to record it. The first one records 70% of

the time, and the second one records 60% of the time. What is the

probability that he gets home after the exam and finds

(a) No copies of his program?

(b) One copy of his program?

(c) Two copies of his program?

Answers:

(a) P(No copies of his program) "= (1-0.70) \times (1-0.60) =0.12"

(b) P(One copy of his program) "= 0.70(1-0.60)+0.60(1-0.70) = 0.46"

(с) P(Two copies of his program) "= 0.70 \times 0.60 = 0.42"